A financial planner wants toinvest 8000 some in stocks earning 15% annually and the rest in bonds earning 6% annually. How much should be invested at each rate to get a return of 930$ annually from the two investments?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

Part I 15.00% per annum ------------- Amount invested =x
Part II 6.00% per annum ------------ Amount invested = y
8000
Interest----- 930.00

Part I 15.00% per annum ---x
Part II 6.00% per annum ---y
Total investment
x + 1 y= 8000 -------------1
Interest on both investments
15.00% x + 6.00% y= 930
Multiply by 100
15 x + 6 y= 93000.00 --------2
Multiply (1) by -15
we get
-15 x -15 y= -120000.00
Add this to (2)
0 x -9 y= -27000
divide by -9
y = 3000 - x)
Part I 15.00% $ 5000
Part II 6.00% $ 3000

CHECK
5000 --------- 15.00% ------- 750.00
3000 ------------- 6.00% ------- 180.00
Total -------------------- 930.00

m.ananth@hotmail.ca

Answer by stanbon(75887)

A financial planner wants to invest 8000 some in stocks earning 15% annually and the rest in bonds earning 6% annually.
How much should be invested at each rate to get a return of 930$ annually from the two investments?
-----------------------
Equations:
Quantity Eq::: f + s = 8000
Interest Eq::: 15f +6s = 93000
---------
Modify for elimination::
6f + 6s = 48000
15f+ 6s = 93000
------
9f = 45000
f = $5000 (amt. invested at 15%)
-------
Solve for "s":
f + s = 8000
s = $3000 (amt. invested at 6%)
---------------------------------
Cheers,
Stan H.
=============